### v1

# Limits to scalable evaluation at the frontier: LLM as Judge won’t beat twice the data

Abstract
High quality annotations are increasingly a bottleneck in the explosively growing machine learning ecosystem. Scalable evaluation methods that avoid costly annotation have therefore become an important research ambition. Many hope to use strong existing models in lieu of costly labels to provide cheap model evaluations. Unfortunately, this method of using models as judges introduces biases, such as self-preferencing, that can distort model comparisons. An emerging family of debiasing tools promises to fix these issues by using a few high quality labels to debias a large number of model judgments. In this paper, we study how far such debiasing methods, in principle, can go. Our main result shows that when the judge is no more accurate than the evaluated model, no debiasing method can decrease the required amount of ground truth labels by more than half. Our result speaks to the severe limitations of the LLM-as-a-judge paradigm at the evaluation frontier where the goal is to assess newly released models that are possibly better than the judge. Through an empirical evaluation, we demonstrate that the sample size savings achievable in practice are even more modest than what our theoretical limit suggests. Along the way, our work provides new observations about debiasing methods for model evaluation, and points out promising avenues for future work.
1 Introduction
As large models continue to advance in their capabilities, it is increasingly challenging for human experts to evaluate newly released models. Expert data annotation is not only slow and costly. Traditional benchmarking also struggles to keep up with rapidly changing model capabilities across an expanding range of tasks. Yet, as models become more powerful, there is hope that models themselves could become powerful tools for scaling up evaluation. An intriguing idea is to use a strong existing model to provide judgments about other models. Such a “model-as-judge” could provide labels to classification instances, compare model outputs, and replace human annotators across a variety of tasks.
Already implemented in practice in numerous instances, the model-as-judge paradigm, however, runs into some roadblocks. When used as judges, models exhibit a range of biases that can skew model comparisons and result in misleading model rankings. Recently proposed debiasing methods, however, promise a compelling way forward. Using a small number of ground truth labels, these methods can potentially debias a large number of model predictions, thus restoring their utility for benchmarking purposes.
Scalable evaluation is most needed—and most challenging—at the evaluation frontier: Newly released models for which we have little intuition as of yet. What makes this case so challenging is that the new model is likely better than the judge in some ways. In this work, we address the question whether debiasing methods together with the model-as-judge paradigm can, in principle, provide an adequate solution to scalable evaluation at the frontier. However, our theory, supported by empirical evaluation, makes a sobering prediction: Whenever the judge model performs worse at its task than the evaluated model, the optimal debiasing method is no better than using twice the ground truth data. This shows that, although there is merit to debiasing, at the evaluation frontier its economic gains are not greater than a factor-two savings in annotation cost.
1.1 Our contributions
In our theoretical work we focus on a standard evaluation setup that encompasses classification, question answering, arena-style comparisons, and safety evaluations. Given an input $x$, a model $m$ produces an output $m(x)$ that receives a binary ground-truth score $s(x,m(x))$. For example, in classification the accuracy score is $s(x,m(x))$ is $1$ if the model’s output $m(x)$ matches the ground-truth label $y(x)$ and 00 otherwise. A judge model provides a binary proxy score $\tilde{s}(x,m(x))$. We imagine that obtaining ground-truth scores is costly and proxy scores are significantly cheaper.
For a fixed model $m$, our goal is to estimate its true score $\operatorname*{\mathbb{E}}s(x,m(x)),$ where the expectation is taken over an input drawn from a distribution. However due to judge bias, simply using cheap proxy scores $\tilde{s}$ can lead to substantially distorted estimates. To motivate the issue, Figure 1 shows how proxy scores can cause highly misleading rankings of model performance.
(a) GPT-4 as judge (84% accuracy).
(b) LLama3-70b as Judge (79% accuracy)
Figure 1: Model ranks on MMLU in a semi-synthetic setting (a) and for the top-10 models on HELM by July 2024 (b). Ranks are heavily perturbed despite high judge accuracy.
Rather than solely relying on proxy scores, we assume that we have a small number $n$ of samples of ground-truth scores $s$, in addition to a larger number $N$ of proxy scores $\tilde{s}$. A debiasing method is any estimator that provides an unbiased estimate of the true score from the sample of available proxy scores and ground truth scores. In particular, debiasing methods guarantee that ranking models by their estimated performance produces correct model rankings with high probability.
We focus on sample size savings in terms of the ground-truth scores $s$ that debiasing methods can achieve by making use of the proxy scores $\tilde{s}$.
  • We first show that the best possible savings are determined by the Pearson correlation between $s$ and $\tilde{s}$ (Theorem 5).
  • Our main result (Theorem 6) then shows that this correlation is small whenever the judge model performs worse than the evaluated model. In that case the best possible outcome is an effective doubling of the number of samples $s$, no matter how many samples of the proxy score $\tilde{s}$ are available (Corollary 7).
  • In experiments on MMLU and MT-Bench, we empirically confirm that sample size savings of more than a factor two are rare: We only observe them when current state-of-the-art models are used to judge substantially weaker models.
In addition, we show that the commonly reported agreement rate between the judge and the ground truth neither meaningfully restricts the bias introduced by the judge model, nor the sample size savings for debiasing methods. Instead, we propose a novel balanced agreement metric that provides meaningful bounds for sample size savings (Theorem 9).
2 Formal setup and motivation
In this section, we introduce the formal model of evaluation that we work with throughout. We then show how judge bias threatens the validity of LLM-based evaluations, even if judge models are quite accurate.
We focus our analysis on aggregated binary evaluations: Given a prompt $x$, a model $m$ receives a binary score $s(x,m(x))$ depending on its output $m(x)$. As both model outputs and evaluation protocols can be non-deterministic, we treat $s(x,m(x))$ as a random variable, even when the prompt $x$ and model output $m(x)$ are fixed. The model $m$ is then evaluated based on its expected score $\operatorname*{\mathbb{E}}s(X,m(X))$. Here, the expectation is taken over the prompt distribution $X\sim\mathcal{D}$, as well as additional randomness in the scores $s(\cdot,m(\cdot))$ introduced by the evaluation protocol.
This setup covers several important benchmarking and evaluation settings:
  • Accuracy in classification and Q&A benchmarks: Here, the prompt $x$ is an instance to be classified or a question to be answered. Letting $m(x)$ denote the model’s output on input $x,$ the accuracy score $s(x,m)$ equals one whenever $m(x)$ equals the label $y(x)$.
  • Arena-style benchmarks: In this setting $x$ is a prompt. We sample the model’s response $m(x)$ and the response $m^{\prime}(x)$ of another randomly selected model $m^{\prime}$. The score $s(x,m)$ indicates whether $m(x)$ was judged to be a better response than $m^{\prime}(x)$.
  • Safety benchmarks: For this, the prompt $x$ is designed to potentially elicit unsafe behaviour. The score $s(x,m)$ indicates whether the model’s response $m(x)$ is safe.
In each of these settings, we can use a language model as a judge model to score models more cheaply than via ground truth evaluation. To do so, we compute a proxy score $\tilde{s}$ with the help of the judge model. We assume that the proxy score is binary as well. In the classification setting, for example, the judge models could provide the labels for evaluation. The proxy score would then be the classification accuracy with respect to the judge model’s labels. In arena comparisons, the LLM could replace human raters. And in safety benchmarks, an LLM could evaluate the safety of model outputs. Although our motivation focuses on language models as the source of proxy scores, our formal work applies to any proxy score, e.g., scores provided by crowdworkers rather than experts.
Precisely modeling the specific relationship between $x$, $s(x,m)$ and $\tilde{s}(x,m)$ is generally infeasible: Even in the simplest case of classification, this would require a precise formal model of not just the judge’s prediction $\tilde{m}(x)$, but also the relationship between $x$ and the correct labels $y(x)$. We therefore treat the scoring functions $s$ and $\tilde{s}$ as black boxes, as is standard, and focus on the joint distribution of the two induced random variables $s(m):=s(X,m)$ and $\tilde{s}(m):=\tilde{s}(X,m)$.
As both scores $s(m)$ and $\tilde{s}(m)$ are binary, we only require three parameters to specify their joint distribution:
$b(m)$ =$=\operatorname*{\mathbb{P}}\{s(m)=1\}$
$p(m)$ =$=\operatorname*{\mathbb{P}}\{\tilde{s}(m)=s(m)\mid s(m)=1\}$
$q(m)$ =$=\operatorname*{\mathbb{P}}\{\tilde{s}(m)=s(m)\mid s(m)=0\}$
Here, the parameter $b(m)$ represents the model’s expected real score $\operatorname*{\mathbb{E}}s(m)$, while $p(m)$ represents the probability that the evaluations $s(m)$ and $\tilde{s}(m)$ match, conditional on the score $s$ being equal to one (i.e., the true positive rate). Similarly, $q(m)$ represents the probability of $s(m)=\tilde{s}(m)$ conditional on $s(m)=0$ (i.e., the true negative rate).
2.1 Judge bias can strongly perturb rankings
The goal is now to understand how to employ the proxy score $\tilde{s}$ to benchmark a set of models $\mathcal{M}$ in a valid way. Following the reasoning in Salaudeen & Hardt (2024), benchmarks have two purposes: quantifying model performance and ranking models. In turn, a useful LLM judge should accurately estimate each model’s performance, as well as induce the same ranking of evaluated models as the ground truth. For the first goal, the main quantity of interest is the judge bias , defined as
$\operatorname*{\mathrm{JB}}(m)\coloneqq\operatorname*{\mathbb{E}}(\tilde{s}(m)-s(m))=\left(1-q(m)\right)\left(1-b(m)\right)-\left(1-p(m)\right)b(m)\,.$
Whenever $\operatorname*{\mathrm{JB}}(m)$ has large magnitude, the proxy score $\tilde{s}$ misrepresents the performance $b(m)$ of model $m$, even with large sample sizes.
When it comes to ranking models, the relationship between judge bias $\operatorname*{\mathrm{JB}}$ and model rankings is more subtle: If the judge bias $\operatorname*{\mathrm{JB}}(m)$ was constant across all models, rankings would be unaffected. However, $\operatorname*{\mathrm{JB}}(m)$ can strongly depend on the evaluated model $m$. If $\operatorname*{\mathrm{JB}}(m)$ is positive for models with low performance, and negative for better performing models, using the score of the judge $\tilde{s}$ can easily perturb model rankings. Our next proposition provides such an example inspired by empirical work. Specifically, real-world classifiers are not only highly similar to each other (Mania et al., 2019), Mania & Sra (2020) also observe that when one classifier is better than another on real data, it tends to be strictly better in a point-wise sense, for the most part. We therefore consider using a classifier $m$ to evaluate a set $\mathcal{M}$ of strictly better classifiers. In that case, using the proxy evaluations $\tilde{s}$ based on $m$ fully reverses the correct model ranking:
Proposition 1.
Consider a binary classifier $\tilde{m}$ and a set of strictly better binary classifiers $\mathcal{M}$ such that $\tilde{m}(x)=y(x)$ implies $m_{i}(x)=y(x)$ for all $m_{i}\in\mathcal{M}$. Let $\operatorname*{\mathbb{E}}s(m_{i})$ represent the accuracy of model $m_{i}$ evaluated on the correct labels, and $\operatorname*{\mathbb{E}}\tilde{s}(m_{i})$ its accuracy evaluated on predictions of model $\tilde{m}$. Then for $m_{i},m_{j}\in\mathcal{M}$, $\operatorname*{\mathbb{E}}s(m_{i})>\operatorname*{\mathbb{E}}s(m_{j})$ implies $\operatorname*{\mathbb{E}}\tilde{s}(m_{i})<\operatorname*{\mathbb{E}}\tilde{s}(m_{j})$.
We prove Proposition 1 in Appendix C.4. An illustration is provided in Figure 1(a). We use GPT-4 labels to evaluate Claude Sonnet, as well as a set of (ficticious) models on MMLU (Hendrycks et al., 2021). The ficticious models are denoted as ”Claude-Sonnet + $x$” and designed to be strictly better than Claude. This is done by changing wrong predictions of Claude into correct ones until accuracy is improved by $x\%$. Figure 1(b) shows a more realistic setting, with LLama-3 70B judging the top-10 MMLU models according to HELM (Liang et al., 2023). While the ranking is not fully reversed, it is strongly perturbed. This effect can also be observed in prior work: For MT-Bench, the ranking of the best models is not consistently preserved when using LLM judges (Zheng et al., 2024).
2.2 High agreement rate is insufficient for evaluation
We note that all judge models from the last section perform well at their task: LLama-3 70B has $79\%$ accuracy on MMLU, while GPT-4 has $84\%$, and the models used in MT-Bench have $85\%$ agreement with expert annotators on average. Our next proposition explains how it is possible for model rankings to be strongly perturbed despite accurate judges. Before stating it, we need to formally define the agreement $\operatorname*{\mathrm{AG}}(m)$ between score $s$ and proxy score $\tilde{s}$ as
$\operatorname*{\mathrm{AG}}(m)\coloneqq\operatorname*{\mathbb{P}}\{s(m)=\tilde{s}(m)\}=b(m)p(m)+\left(1-b(m)\right)q(m).$
In the classification setting, the judge model’s accuracy lower bounds the agreement $\operatorname*{\mathrm{AG}}(m)$: Clearly $s(x,m)=\tilde{s}(x,m)$ whenever the judge model $\tilde{m}$ is correct. But when there are more than two classes, we also get $s(x,m)=\tilde{s}(x,m)=0$ whenever the true label $y(x)$ and the predictions $m(x)$ and $\tilde{m}(x)$ are all different from each other. In particular, this highlights that the agreement $\operatorname*{\mathrm{AG}}(m)$ can indeed depend on the evaluated model $m$, rather than just the judge $\tilde{m}$. For binary classification, however, the agreement rate equals the judge model’s accuracy and is thus constant across evaluated models. Our next proposition now relates judge bias to agreement:
Proposition 2.
Fix any model score $b(m)$. Then for any $r$ with $b(m)\leq r,$ there are values of $q(m)$ and $p(m)$, such that $\operatorname*{\mathrm{AG}}(m)=r$ and we obtain positive judge bias $\operatorname*{\mathrm{JB}}(m)=1-r$. Similarly, for any $r$ with $1-b(m)\leq r,$ there are values of $q(m)$ and $p(m)$, such that $\operatorname*{\mathrm{AG}}(m)=r$ and we obtain negative judge bias $\operatorname*{\mathrm{JB}}(m)=r-1$.
Proposition 2 shows that for agreement $\operatorname*{\mathrm{AG}}(m)=r$, there can be a judge bias of $1-r$ in either direction. This means that even if $\operatorname*{\mathrm{AG}}(m)=r$ was constant across models $m$, we could only reliably rank models for which the true score $\operatorname*{\mathbb{E}}s(m)$ differs by more than $2(1-r)$. As it is common for state-of-the-art models to differ in performance by only low single digit percentages, without further assumptions an agreement rate above $99\%$ (for each evaluated model!) would be required to ensure (asymptotically) correct rankings. For a proof, see Appendix C.5.
3 Debiasing LLM Judgments
In the previous section, we demonstrated how biased LLM judges can lead to strongly perturbed model rankings. This motivates the need for debiasing methods. We begin by discussing a simple sufficient condition to guarantee asymptotically correct rankings: Ensure that our estimates of $\operatorname*{\mathbb{E}}s(m)$ are unbiased for all models $m$. Statistical bias correction methods can make use of a large number of proxy scores and a smaller number of gold standard labels to construct an unbiased estimator. We discuss one such method in detail, and show that it is essentially optimal in terms of estimator variance. Afterwards we prove our main result, i.e., that LLM judges offer limited benefits for evaluating state-of-the-art models. We conclude the section with additional upper and lower bounds on the benefits of judge labels based on a novel balanced agreement metric.
3.1 Uniformly controlling judge bias ensures correct rankings
We begin with a simple and natural sufficient condition for correct rankings. This is that all model evaluations are approximately unbiased. The next proposition shows that this, in fact, entails correct rankings in expectation and hence also with a sufficient amount of data.
Proposition 3.
Let $M$ be a finite set of models such that for any $m,m^{\prime}\in\mathcal{M}$
$|\operatorname*{\mathbb{E}}(s(m))-\operatorname*{\mathbb{E}}(s(m^{\prime}))|\geq\epsilon.$
Let $\hat{\theta}$ be an estimator for $\operatorname*{\mathbb{E}}s(m)$ such that $\operatorname*{\mathrm{Var}}\hat{\theta}(m)$ converges to zero in the dataset size. Then if $|\operatorname*{\mathbb{E}}\hat{\theta}(m)-\operatorname*{\mathbb{E}}(s(m))|<\frac{\epsilon}{2}$ for all models $m$, using $\hat{\theta}$ to rank models yields the correct ranking with high probability.
Proof.
By construction, $\operatorname*{\mathbb{E}}\hat{\theta}(m)$ induces the same model rankings as $\operatorname*{\mathbb{E}}s(m)$. Chebyshev’s inequality and a union bound imply correct rankings as $\operatorname*{\mathrm{Var}}\hat{\theta}(m)$ goes to zero. ∎
In particular, this implies that if $|\operatorname*{\mathrm{JB}}(m)|\leq\frac{\epsilon}{2}$ for all models $m\in\mathcal{M}$, a sufficiently large amount of proxy samples $\tilde{s}$ will give us the correct rankings with high probability. However, it is important for the condition to hold for all models. Judge bias might be low averaged over models, even if the judge overestimates the performance of some models and underestimates the performance of others.
3.2 Background on Prediction Powered Inference
While Proposition 3 suggests an avenue to obtain correct model rankings when using a proxy score, actually ensuring a small judge bias $|\operatorname*{\mathrm{JB}}(m)|$ for all evaluated models would require near-perfect judge models. Instead, we switch our attention to recently proposed statistical debiasing methods that make use of a small amount of ground truth data to debias a given proxy score.
Specifically, we follow the Prediction Powered Inference (PPI) framework (Angelopoulos et al., 2023a) and an for our case equivalent method by Chaganty et al. (2018): Alongside a large iid sample of model judgments $\tilde{s}_{i}(m)=\tilde{s}(x_{i},m)$ for integer $i\in[1,N+n]$, we assume access to the corresponding ground truth label $s_{j}(m)=s(x_{j},m)$ for a small subset of integer $j\in[1,n]$. With this, we can construct an unbiased estimator $\hat{\theta}^{PP}$ for $\operatorname*{\mathbb{E}}s(m)$ by using the small parallel sample to estimate $\operatorname*{\mathrm{JB}}(m)$, and subtracting it from our estimate of $\operatorname*{\mathbb{E}}\tilde{s}(m)$. The PPI estimator thus equals
$\hat{\theta}^{PP}(x)=\frac{1}{N}\sum_{i=n+1}^{N+n}\tilde{s}_{i}(m)+\frac{1}{n}\sum_{j=1}^{n}\left(s_{j}(m)-\tilde{s}_{j}(m)\right).$
Clearly, the PPI estimator is unbiased
$\operatorname*{\mathbb{E}}\hat{\theta}^{PP}=\operatorname*{\mathbb{E}}\tilde{s}(m)+\operatorname*{\mathbb{E}}s(m)-\operatorname*{\mathbb{E}}\tilde{s}(m)=\operatorname*{\mathbb{E}}s(m),$
and has variance
$\operatorname*{\mathrm{Var}}\hat{\theta}^{PP}=\frac{1}{N}\operatorname*{\mathrm{Var}}\tilde{s}(m)+\frac{1}{n}\operatorname*{\mathrm{Var}}(\tilde{s}(m)-s(m)).$
Whenever $N\gg n$, the first term on the right hand side is small. Moreover, when $s(m)$ and $\tilde{s}(m)$ are strongly correlated, the second term is small. Assuming both, we have $\operatorname*{\mathrm{Var}}\hat{\theta}^{PP}<\operatorname*{\mathrm{Var}}\hat{\theta}^{GT}$, where $\hat{\theta}^{GT}$ is the ground-truth sample average estimator $\hat{\theta}^{GT}=\frac{1}{n}\sum_{j=1}^{n}s_{j}(m).$
However if the correlation between $s(m)$ and $\tilde{s}(m)$ is small, the inequality is reversed. This issue can be fixed by interpolating between the PPI estimator and the classical estimator (Chaganty et al., 2018; Angelopoulos et al., 2023b), setting
$\hat{\theta}^{PP}_{\lambda}=\lambda\hat{\theta}^{PP}+(1-\lambda)\hat{\theta}^{GT}.$
As a linear combination of unbiased estimators this remains unbiased, but we can now optimize over $\lambda$ to minimize estimator variance. We obtain the optimum value
$\lambda^{*}=\frac{\operatorname*{\mathrm{Cov}}(s(m),\tilde{s}(m))}{\left(1+\frac{n}{N}\right)\operatorname*{\mathrm{Var}}\tilde{s}(m)}.$
The estimator $\hat{\theta}^{PP}_{\lambda^{*}}$never increases variance compared to the classical ground truth estimator:
$\operatorname*{\mathrm{Var}}\hat{\theta}^{PP}_{\lambda^{*}}=\frac{1}{n}\operatorname*{\mathrm{Var}}s(m)-\frac{1}{n+\frac{n^{2}}{N}}\frac{\operatorname*{\mathrm{Cov}}(s(m),\tilde{s}(m))^{2}}{\operatorname*{\mathrm{Var}}(\tilde{s}(m))}\leq\frac{1}{n}\operatorname*{\mathrm{Var}}s(m)=\operatorname*{\mathrm{Var}}\hat{\theta}^{GT}.$
As the variance of $\hat{\theta}^{GT}$ scales as $\Theta(1/n)$, using an estimator $\hat{\theta}$ with $r=\frac{\operatorname*{\mathrm{Var}}\hat{\theta}}{\operatorname*{\mathrm{Var}}\hat{\theta}^{GT}}$ is equivalent to increasing the ground-truth sample size by a factor of $\tau(\hat{\theta})\coloneqq\frac{1}{r}.$ We call $\tau(\hat{\theta})$ the sample efficiency factor of the estimator $\hat{\theta}$. For fixed values of $n$ and $N$, the relative variance reduction of $\hat{\theta}^{PP}_{\lambda^{*}}$ compared to $\hat{\theta}^{GT}$ is determined by the squared Pearson correlation coefficient $\rho$ between the ground truth labels $s$ and model judgements $\tilde{s}$:
$\frac{\operatorname*{\mathrm{Var}}\hat{\theta}^{PP}_{\lambda^{*}}}{\operatorname*{\mathrm{Var}}\hat{\theta}^{GT}}=1-\frac{1}{1+\frac{n}{N}}\rho(s(m),\tilde{s}(m))^{2}.$
Therefore, we have
$\tau(\hat{\theta}^{PP}_{\lambda^{*}})=\frac{1}{1-\frac{1}{1+\frac{n}{N}}\rho(s(m),\tilde{s}(m))^{2}}\geq 1$
The sample efficiency factor $\tau(\hat{\theta}^{PP}_{\lambda^{*}})$ is large if the squared correlation between true and proxy scores is high. Conversely, if the squared correlation is small, so is the sample efficiency factor.
3.3 PPI is near-optimal for binary evaluations
As the correlation $\rho$ is central to the sample efficiency factor $\tau$, we assume $b(m),p(m),q(m)\in(0,1)$ for the remainder of the text to exclude extreme cases in which $\rho(s(m),\tilde{s}(m))$ is undefined. It is now easy to upper bound the sample efficiency factor:
Proposition 4.
The sample efficiency factor for the PPI estimator is upper bounded by
$\tau(\hat{\theta}^{PP}_{\lambda^{*}})\leq\frac{1}{1-\rho(s(m),\tilde{s}(m))^{2}},$
where
$\rho(s(m),\tilde{s}(m))^{2}=\frac{b(m)}{(1-b(m))}\frac{(p(m)-\operatorname*{\mathbb{E}}\tilde{s}(m))^{2}}{\operatorname*{\mathbb{E}}\tilde{s}(m)(1-\operatorname*{\mathbb{E}}\tilde{s}(m))}.$
The proof can be found in Appendix C.6.
This maximal variance reduction is only attained asymptotically: First, it requires unlimited proxy data, i.e. $N\to\infty$. Second, it assumes that the optimal $\lambda^{*}$ is known. In practice, $\lambda^{*}$ has to be estimated from data. This introduces additional variance and comes at the price of a bias of order $O(1/n)$ (Chaganty et al., 2018), if $\lambda^{*}$ is fit on the same data, or reduced sample size, if some data is set aside for estimating $\lambda^{*}$.
It is also worth noting that unless we have perfect correlation $\rho^{2}=1$, the sample efficiency factor $\tau(\hat{\theta}^{PP}_{\lambda^{*}})$ is finite. For any correlation bounded away from $1$, the proxy samples can only provide a constant factor improvement. Scaling them up without also scaling the number of ground truth samples has limited benefits. This raises the question whether we can find an unbiased estimator with lower variance. For black-box estimators that do not attempt to model the distribution of $(s(m),\tilde{s}(m))$ conditional on the prompt $x$, we show that the answer is no:
Theorem 5.
Any unbiased estimator $\hat{\theta}$ for $\operatorname*{\mathbb{E}}s(m)$ that observes $n$ joint samples $(s(m),\tilde{s}(m))$ and $N$ independent proxy samples $\tilde{s}(m)$ fulfills the variance lower bound
$\operatorname*{\mathrm{Var}}\hat{\theta}\geq\operatorname*{\mathrm{Var}}\hat{\theta}^{PP}_{\lambda^{*}}\,.$
This means that the sample efficiency factor of $\hat{\theta}$ is bounded by
$\tau(\hat{\theta})\leq\tau(\hat{\theta}^{PP}_{\lambda^{*}})\leq\frac{1}{1-\rho(s(m),\tilde{s}(m))^{2}}\,.$
Theorem 5 follows from an application of the Cramér-Rao bound, combined with extensive algebraic manipulation that was assisted by a computer algebra system. Appendix C.1 has the details.
3.4 Limited gains for evaluating state-of-the-art models
With Theorem 5 at hand, we can analyze the effective sample size gains $\tau$ that unbiased estimators can achieve by making use of model judgments $\tilde{s}$. In this subsection, we focus on the evaluation of frontier models that outperform older models, including the models used to judge them. We use the evaluated model’s score $b(m)$ and the judge’s agreement rate $\operatorname*{\mathrm{AG}}(m)$ as indicators for the respective model’s capabilities. We focus on tasks for which performing and judging the task are of similar difficulty, such that these indicators are comparable. As a prime example, for binary classification, $\operatorname*{\mathrm{AG}}(m)$ equals the judge’s accuracy, while $b(m)$ equals the evaluated model’s accuracy.
We thus capture the evaluation of frontier models with the assumption that the evaluated model’s score $b(m)$ is larger than the judge’s agreement rate $\operatorname*{\mathrm{AG}}(m)$. The next theorem shows that the squared correlation $\rho(s(m),\tilde{s}(m))^{2}$ is at most one half in this setting.
Theorem 6.
Assume $0.5\leq\operatorname*{\mathrm{AG}}(m)\leq b(m)$. Then,
$\rho(s(m),\tilde{s}(m))^{2}\leq 0.5.$
We prove Theorem 6 in Appendix C.2. Combining Theorem 6 with Theorem 5 allows us to bound the sample efficiency gains that any unbiased estimator can achieve for evaluating frontier models:
Corollary 7.
Assume $0.5\leq\operatorname*{\mathrm{AG}}(m)\leq b(m)$ and let $\Theta$ be the set of all unbiased estimators for $\operatorname*{\mathbb{E}}s(m)$ that observe $n$ joint samples $(s(m),\tilde{s}(m))$ and $N$ independent proxy samples $\tilde{s}(m)$. Then all estimators in $\Theta$ have a bounded sample efficiency factor:
$\tau_{\max}\coloneqq\max_{\hat{\theta}\in\Theta}\tau(\hat{\theta})\leq 2.$
Corollary 7 implies that when evaluating state-of-the-art models, the best we can expect from using LLM judges is a factor-two improvement in sample efficiency. This is unless the judge’s task is substantially easier than the evaluated model’s task.
3.5 High agreement rate is insufficient for sample efficiency gains
Theorem 6 holds for arbitrarily high levels of agreement. Therefore, the theorem shows that high agreement rates are not just insufficient for avoiding judge bias—they’re also insufficient for obtaining a meaningful sample efficiency factor $\tau$. In this subsection, we go a step further and show that the agreement $\operatorname*{\mathrm{AG}}(m)$ does not provide any lower bound on the sample efficiency factor $\tau$. Addressing this issue, we then propose a modified balanced agreement score and show that it can provide meaningful upper and lower bounds on $\tau$.
Our first proposition shows that agreement does not meaningfully constrain the squared correlation $\rho^{2}$, even when $\operatorname*{\mathrm{JB}}(m)$ is small. Correspondingly, in some cases access to proxy scores $\tilde{s}$ does not improve sample efficiency at all, even if $\tilde{s}$ turns out to be unbiased.
Proposition 8.
For any agreement rate $10.5\leq\operatorname*{\mathrm{AG}}(m)<$ 1, there exist values of $b(m),p(m),q(m)\in(0,1)$ such that
$\operatorname*{\mathrm{JB}}(m)=\rho(s(m),\tilde{s}(m))^{2}=0.$
Proof.
We drop the dependence of our parameters on $m$ and set $q=1-b$ and $p=b$. Then $\operatorname*{\mathrm{JB}}=b(1-b)-(1-b)b=0.$ At the same time, as there is no judge bias, $\operatorname*{\mathbb{E}}\tilde{s}(m)=b=p$, such that $\rho(s(m),\tilde{s}(m))^{2}=0$. Lastly $\operatorname*{\mathrm{AG}}=b^{2}+(1-b)^{2}$, which can attain any value between $0.5$ and $1$. ∎
In particular, the maximal sample efficiency factor $\tau_{\max}$ equals one in that case, such that despite access to $\tilde{s}$, no unbiased estimator has less variance than $\hat{\theta}^{GT}$. It might be surprising that there are cases in which we cannot make productive use of the proxy evaluations $\tilde{s}$ despite zero judge bias. However, without further assumptions, $\operatorname*{\mathrm{JB}}(m)$ can vary wildly between models, so that we cannot skip estimating it anew for each model. This is true even in cases where $\operatorname*{\mathrm{JB}}(m)$ turns out to be close to zero, post-hoc. Now, the estimator does not ”know” that $\tilde{s}$ has zero judge bias, but rather has to estimate $\operatorname*{\mathrm{JB}}(m)$ from joint observations $(s,\tilde{s})$. In the configuration from the proposition, estimating the judge bias $\operatorname*{\mathrm{JB}}(m)$ turns out to be as hard as directly estimating the real score $\operatorname*{\mathbb{E}}s(m)$.
3.6 Balanced agreement rate as an alternative
Intuitively, agreement fails to constrain the squared correlation $\rho^{2}$ because predictors without any signal, such as constant predictors, can achieve high agreement. This is analogous to how accuracy can be misleading for imbalanced classification problems. In classification, this is solved by focusing on balanced accuracy instead. By analogy, we introduce the balanced agreement of $\tilde{s}$ with $s$ as
$\operatorname*{\mathrm{BA}}(m)\coloneqq\frac{q(m)+p(m)}{2}.$
The next theorem confirms that controlling $\operatorname*{\mathrm{BA}}(m)$ yields meaningful lower bounds on the squared correlation.
Theorem 9.
For any value of $\operatorname*{\mathrm{BA}}(m)$, we have:
$4b(m)(1-b(m))(2\operatorname*{\mathrm{BA}}(m)-1)^{2}\leq\rho(s(m),\tilde{s}(m))^{2}\leq|2\operatorname*{\mathrm{BA}}(m)-1|\,.$
Theorem 9 is proven in Appendix C.3. It also provides another upper bound on sample efficiency gains from access to $\tilde{s}$: Whenever $\operatorname*{\mathrm{BA}}(m)$ is close to half, access to $\tilde{s}$ can not improve sample efficiency by much and the sample efficiency factor $\tau$ is close to one. As a corollary, we show that whenever the proxy has reasonable balanced accuracy, the minimum of the true positive rate $p$ and true negative rate $q$ upper bounds the squared correlation $\rho^{2}$:
Corollary 10.
Whenever $\operatorname*{\mathrm{BA}}(m)\geq 0.5$, we have
$\rho(s(m),\tilde{s}(m))^{2}\leq\min\{p(m),q(m)\}.$
Proof.
Given $\operatorname*{\mathrm{BA}}(m)\geq 0.5$, $|2\operatorname*{\mathrm{BA}}(m)-1|$ is maximized at the maximal value of $\operatorname*{\mathrm{BA}}(m)$. But with $\min\{p(m),q(m)\}\leq r$, that is at $\frac{r+1}{2}$. Inserting into Theorem 9 yields the result. ∎
4 Experiments
Our experiments focus on two settings: Multiple-choice question answering and chatbot evaluation. For the first, we evaluate the accuracy of models on MMLU (Hendrycks et al., 2021), a multiple-choice question answering benchmark that aims to measure the world knowledge of language models. We aggregate all subtask test sets of MMLU to obtain a single test set $T$, and measure the performance of a model $m$ on MMLU in terms of accuracy $\frac{1}{|T|}\sum_{x}\mathbb{I}(m(x)=y(x))$. Here $\mathbb{I}$ is the indicator function, $m(x)$ is the model’s answer and $y(x)$ is the correct answer. For model-based evaluations using model $\tilde{m}$ as a judge, we simply replace the correct answer $y(x)$ by the judge’s prediction $\tilde{m}(x)$ and calculate accuracy based on the modified answers.
Figure 2: Best possible sample efficiency factor $\tau_{\max}$ using different judges (colors) evaluating different models (x-ticks) on MMLU. Error bars show $90\%$ confidence intervals.
In the second setting, we evaluate model performance on MT-Bench (Zheng et al., 2024), an arena-style benchmark designed to evaluate the ability of models to follow instructions. The benchmark comes with results for six models. For each pair of these models $m$ and $m^{\prime}$, both models are queried with the same prompt $x$. The evaluator then decides, which model’s response was better. For each model $m$, we calculate its win rate averaged over all triples of the form $(m,m^{\prime},x)$, for other models in the benchmark $m^{\prime}$ and prompts $x$. For more details, consider Appendix B.
Figure 3: Best possible sample efficiency factor $\tau_{\max}$ using GPT-4-as-a-Judge, evaluating different models on MT-Bench. Error bars show $90\%$ confidence intervals.
Figures 2 and 3 show the optimal sample efficiency factor $\tau_{\max}$ on MMLU and MT-Bench respectively. The sample efficiency factor $\tau_{\max}$ consistently stays below the value of two implied by Corollary 7, even when slightly stronger models like GPT-4 are used to evaluate weaker models like LLama3-70B. We only get a larger sample efficiency factor $\tau_{\max}$ when current flagship models are used to judge the significantly worse LLama2-7b on MMLU. Notably, $\tau_{\max}$ is often substantially smaller than two, especially on MMLU. This highlights that Corollary 7 is a best-case upper bound rather than a guarantee that sample efficiency factors close to two are actually achievable in practice.
Going beyond binary evaluations.
We also experiment with incorporating model uncertainty into the proxy score on MMLU. We define the proxy score as the (renormalized) probability the judge model $\tilde{m}$ assigned to model $m$’s answer $\tilde{s}(m)=\operatorname*{\mathbb{P}}_{\tilde{m}(x)}\{m(x)\}$. While this non-binary score $\tilde{s}$ breaks the assumptions of Theorem 5, we expect PPI to remain practically optimal. Figure 4 shows that for LLama3.1-405B as judge, this approach increases the sample efficiency of PPI. However, when evaluating the strongest models, the sample efficiency factor $\tau(\hat{\theta}_{\lambda^{*}}^{PP})$ still remains below two.
Figure 4: Data efficiency factor $\tau(\hat{\theta}_{\lambda^{*}}^{PP})$ in the $N\to\infty$ limit, using LLama3.1-405B-as-a-Judge on MMLU, with accuracy scoring vs soft accuracy scoring. Error bars show $90\%$ confidence intervals.
5 Related work
LLM-as-a-Judge.
With the success of large transformer-based language models, the idea of using model predictions to provide feedback for another model has become very popular. Initially, specialized fine-tuned models were used to provide training signals for a second model (Ouyang et al., 2022; Bai et al., 2022; Dorner et al., 2023). The approach was quickly expanded to using general purpose LLMs like GPT-4 (Achiam et al., 2023), not just for providing training signal but also for evaluating other models. This new paradigm, dubbed LLM-as-a-Judge, has been applied to evaluate a variety of model capabilities (Yu et al., 2023; Chiang & Lee, 2023; Fu et al., 2023; Li et al., 2024; Weyssow et al., 2024; Raju et al., 2024; Vu et al., 2024; Kumar et al., 2024). In some cases, not just the ratings, but also the prompts given to evaluated models are designed by LLMs (Bai et al., 2024). LLM-as-a-Judge is often paired with another emerging paradigm for model evaluation: As generative tasks often do not have a single correct answer, models are evaluated in so-called arena benchmarks (Chiang et al., 2024), where different models’ responses to the same prompts are ranked to determine the best model. Beyond evaluating models, LLM judges have also been employed to evaluate red-teaming (Mazeika et al., 2024) and jailbreaks (Souly et al., 2024). The use of LLM judges rather than experts or crowd workers is often justified by high agreement rates (Gilardi et al., 2023; Zheng et al., 2024) between both types of judges. However, Thakur et al. (2024) find that seemingly high agreement does not necessarily imply accurate judge scores. In this work, we provide a theoretical justification for that finding and show that agreement is not necessarily a good indicator of judge quality, even when judge scores are debiased via PPI.
Bias in LLM judges.
LLM judges can be biased in numerous ways, making their evaluations unreliable: Their outputs often correlate poorly with expert annotations (Bavaresco et al., 2024; Koo et al., 2023). Models are known to rate their own outputs more favorable than the outputs of other models (Liu et al., 2023; Panickssery et al., 2024), prefer longer outputs and outputs containing lists, regardless of quality (Dubois et al., 2024b; Wei et al., 2024). They also exhibit choice-order bias (Dominguez-Olmedo et al., 2023; Wang et al., 2023; Shi et al., 2024) as well as a variety of other biases (Koo et al., 2023). While correcting for these known biases (Dubois et al., 2024a; Zheng et al., 2024) is an important first step, manually enumerating and correcting for all judge biases appears infeasible. As one potential alternative, Jung et al. (2024) propose to combat judge bias by having judges abstain based on their confidence. However, if abstention correlates with the evaluated model’s performance, this approach can introduce its own biases to model evaluations.
Debiasing methods.
Usually, the biases listed above are found by identifying patterns in how LLM judgments deviate from (a smaller set of) ground truth labels. Instead of trying to identify and fix specific biases, another line of work uses ground truth labels to directly estimate the bias an LLM judge introduces and correct for it. This approach was already suggested by Chaganty et al. (2018) for debiasing classic automated NLP metrics like BLEU (Papineni et al., 2002) and ROUGE (Lin & Och, 2004). The authors find that their method, which is essentially equivalent to PPI, only improved data efficiency by around $10\%$ using 2018’s automated metrics. In addition, they show that their method has optimal worst-case variance. In comparison to their worst-case result, our Theorem 5 shows that $\hat{\theta}^{PP}_{\lambda^{*}}$is always optimal in our binary evaluation setting.
For modern LLM-based metrics, using PPI has been suggested by Boyeau et al. (2024) and Chatzi et al. (2024), and applied as part of more complicated evaluation pipelines (Saad-Falcon et al., 2023; Tyser et al., 2024). Furthermore, Fisch et al. (2024) combine PPI with stratified sampling for model evaluation. These works consistently show that PPI improves efficiency in terms of ground truth labels, but gains in effective sample size rarely exceed $50\%$ and are always below $100\%$. Our upper bounds on the sample efficiency factor $\tau_{\max}$ show that gains larger than that are indeed unlikely, especially when evaluating state-of-the-art models.
6 Discussion
Our results show that for evaluating frontier models, LLM judges might fall short of the promise of largely replacing expert labelers: While doubling the effective sample size can be useful for practitioners, the order of magnitude of required ground truth labels remains the same with and without access to LLM judges. That said, there are ways to circumvent our negative results:
First, our results pertain to uniform sampling rather than more sophisticated sampling strategies. Approaches like stratified PPI (Fisch et al., 2024) might be able to obtain a somewhat better sample efficiency factor, for example by breaking the assumptions of Theorem 6 per-stratum. Second, Theorem 6 only holds if the proxy $\tilde{s}$ is less capable at predicting $s$ than the evaluated model $m$ is at its task. Thus, for tasks in which evaluation is substantially easier than the task itself, or tasks for which data to train a specialized evaluator is abundant, gains in sample efficiency of more than a factor two are possible even when evaluating models at the frontier. In addition, weaker models might find their way into many applications due to cost saving reasons, such that efficient evaluations of these models are still valuable. That said, Theorem 5 still applies per-stratum, as well as for weaker models. Correspondingly, variance reduction is still limited by the squared correlation $\rho(s(m),\tilde{s}(m))^{2}$.
Finally, we would like to reiterate that our results do not only apply to LLM judges, but any form of biased evaluators including (poorly instructed) crowdworkers. Whenever the crowdworker majority does not consistently agree with the desired ground truth labels, bias correction is required for valid evaluation results. Similarly, if the crowdworker majority vote is less accurate than the evaluated model, Theorem 6 implies that crowdworker labels are of low value compared to ground truth labels.
7 Acknowledgements
We would like to thank Anastasios Angelopoulos, Amin Charusaie, Yatong Chen, André Cruz and Guanhua Zhang for helpful discussions and/or feedback on draft versions of this work. Florian Dorner and Vivian Nastl are grateful for financial support from the Max Planck ETH Center for Learning Systems (CLS).
Appendix A Additional theoretical results
A.1 Sample efficiency factor for strictly better classifiers
In the case of strictly better classifiers discussed in Proposition 1, we can provide an even stronger upper bound: When using a classifier $\tilde{m}$ with accuracy $x>0.5$ to evaluate a strictly better classifier $m$ with accuracy $x+\delta$, we get $q(m)=0$, $b(m)=x+\delta$ as $p(m)=\frac{x}{x+\delta}$. In that case we can only reduce variance by a factor of $\frac{1}{9}$, or equivalently improve sample efficiency by a factor of $1.125$:
Proposition 11.
Fix $q(m)=0$ and assume $b(m)=x+\delta$ as well as $p(m)=\frac{x}{x+\delta}$ for $x>0.5$, $\delta>0$ and $x+\delta\leq 1$. Then
$\rho(s(m),\tilde{s}(m))^{2}\leq\frac{1}{9}.$
Correspondingly, the sample efficiency factor is bounded by $\tau_{\max}\leq 1.125$.
Proof.
Fixing $q(m)=0$, we get
$\rho(s(m),\tilde{s}(m))^{2}=\frac{\left(\frac{x}{x+\delta}-1\right)\left(\delta+x-1\right)}{1-\delta}.$
Taking the $\delta$ derivative yields
$\frac{d}{d\delta}\rho(s(m),\tilde{s}(m))^{2}=-\frac{x(2\delta+x-1)}{(\delta-1)^{2}(\delta+x)^{2}}.$
This is zero precisely when $\delta=\frac{1-x}{2}$, and negative for $\delta$ larger than that. Correspondingly, $\delta=\frac{1-x}{2}$ is the global maximum. Inserting this back, we obtain
$\rho(s(m),\tilde{s}(m))^{2}\leq\frac{(x-1)^{2}}{(x+1)^{2}}$
with negative derivative $\frac{d}{dx}\frac{(x-1)^{2}}{(x+1)^{2}}=\frac{4(x-1)}{(x+1)^{3}}$. Correspondingly, the upper bound is maximized at the smallest possible $x=0.5$, where it equals $\frac{0.5^{2}}{1.5^{2}}=\frac{1}{9}$. ∎
A.2 Estimating the gap
If ranking rather than accurately estimating model performance is the main goal, we need to estimate the difference $\operatorname*{\mathbb{E}}(s(m))-\operatorname*{\mathbb{E}}(s(m^{\prime}))$ for model pairs $(m,m^{\prime})$. Because of potential cancelations, the difference of the optimal estimators for both terms $\hat{\theta}^{PP}_{\lambda^{*}(m)}(m)-\hat{\theta}^{PP}_{\lambda^{*}(m^{\prime})}(m^{\prime})$ is not necessarily the optimal estimator for the difference. Correspondingly, further improvements can be made by jointly optimizing $\lambda$ and $\lambda^{\prime}$ to minimize the variance of $\hat{\theta}^{PP}_{\lambda}(m)-\hat{\theta}^{PP}_{\lambda^{\prime}}(m^{\prime})$. Solving the corresponding optimization problem yields:
$\lambda^{*}(m)$ $=\frac{\operatorname*{\mathrm{Cov}}(s(m^{\prime}),\tilde{s}(m^{\prime}))\operatorname*{\mathrm{Cov}}(\tilde{s}(m),\tilde{s}(m^{\prime}))-\operatorname*{\mathrm{Cov}}(s(m),\tilde{s}(m^{\prime}))\operatorname*{\mathrm{Cov}}(\tilde{s}(m),\tilde{s}(m^{\prime}))}{(1+\frac{n}{N})(\operatorname*{\mathrm{Var}}(\tilde{s}(m^{\prime}))\operatorname*{\mathrm{Var}}(\tilde{s}(m))-\operatorname*{\mathrm{Cov}}(\tilde{s}(m),\tilde{s}(m^{\prime})))}$
$+\frac{\operatorname*{\mathrm{Cov}}(s(m),\tilde{s}(m))\operatorname*{\mathrm{Var}}(\tilde{s}(m^{\prime}))-\operatorname*{\mathrm{Cov}}(s(m^{\prime}),\tilde{s}(m))\operatorname*{\mathrm{Var}}(\tilde{s}(m^{\prime}))}{(1+\frac{n}{N})(\operatorname*{\mathrm{Var}}(\tilde{s}(m^{\prime}))\operatorname*{\mathrm{Var}}(\tilde{s}(m))-\operatorname*{\mathrm{Cov}}(\tilde{s}(m),\tilde{s}(m^{\prime})))}$
$\lambda^{*}(m^{\prime})$ $=\frac{\operatorname*{\mathrm{Cov}}(s(m),\tilde{s}(m))\operatorname*{\mathrm{Cov}}(\tilde{s}(m),\tilde{s}(m^{\prime}))-\operatorname*{\mathrm{Cov}}(s(m^{\prime}),\tilde{s}(m))\operatorname*{\mathrm{Cov}}(\tilde{s}(m),\tilde{s}(m^{\prime}))}{(1+\frac{n}{N})(\operatorname*{\mathrm{Var}}(\tilde{s}(m^{\prime}))\operatorname*{\mathrm{Var}}(\tilde{s}(m))-\operatorname*{\mathrm{Cov}}(\tilde{s}(m),\tilde{s}(m^{\prime})))}$
$+\frac{\operatorname*{\mathrm{Cov}}(s(m^{\prime}),\tilde{s}(m^{\prime}))\operatorname*{\mathrm{Var}}(\tilde{s}(m))-\operatorname*{\mathrm{Cov}}(s(m),\tilde{s}(m^{\prime}))\operatorname*{\mathrm{Var}}(\tilde{s}(m))}{(1+\frac{n}{N})(\operatorname*{\mathrm{Var}}(\tilde{s}(m^{\prime}))\operatorname*{\mathrm{Var}}(\tilde{s}(m))-\operatorname*{\mathrm{Cov}}(\tilde{s}(m),\tilde{s}(m^{\prime})))}$
However, using different values of $\lambda$ for different comparisons means that there are multiple different score estimates for model $m$ such that the resulting comparisons might not yield a valid transitive ranking.
Appendix B Additional details on experiments
For MMLU, we obtain most model predictions from the HELM (Liang et al., 2023) leaderboard and focus on the top 10 models [^1] Cutoff date: 23.07.2024. As HELM does not document model uncertainty, and results for LLama3.1-405B were initially not available, we evaluated LLama3.1-405B in bf16 using 8 A100 GPUs and the accelerate library for offloading, ourselves. We use the prompting format from HELM and extract the predicted probabilities $\tilde{p}(y)$ corresponding to the four tokens $y\in Y$ that represent the answer options. For our uncertainty experiment, we renormalize them, setting $p(y)=\frac{\tilde{p}(y)}{\sum_{y^{\prime}\in Y}\tilde{p}(y^{\prime})}$.
For MT-bench, we use the results for six models released by the benchmark’s authors Zheng et al. (2024). For each triple consisting of two different models $m$, $m^{\prime}$ and a prompt $x$ there is a varying amount of human expert judgments, as well as a judgment by GPT-4. We aggregate the expert judgments using a majority vote, and discard all triples for which the expert’s tied, or GPT-4’s judgment amounts to a tie. We then calculate the win-rate for model $m$ by averaging over all remaining triples that include $m$. Note, that this is slightly different from the evaluation in the MT-Bench paper, where win-rates are calculated per model pair, and then averaged over models. MT-Bench also includes a follow-up prompt, with separate judgments for the first and second response. For simplicity, we focus on the first answer.
High agreement is not necessary for ranking.
Proposition 2 indicates that a high agreement rate $\operatorname*{\mathrm{AG}}(m)$ is not sufficient for the proxy score $\tilde{s}$ to yield similar model rankings as the real score $s$. However, it turns out that high agreement is also not necessary for stable rankings: For binary classification with judge errors that are independent of the performance of different models, a sufficiently large sample size guarantees correct rankings as long as the error rate is below $50\%$ (Dorner & Hardt, 2024). We demonstrate this phenomenon in Figure 5. In that experiment, the judge reproduces the correct label $60\%$ of the time, but picks a random wrong label otherwise. Despite this judge’s low agreement with the correct labels, model rankings are preserved near-perfectly.
Figure 5: Model rankings on MMLU using a 60% accurate judge with random errors. Ranks are preserved despite agreement substantially below LLM judges.
Appendix C Proofs
C.1 Theorem 5
We use a multivariate version of the Cramér-Rao bound, as presented in Lehmann & Casella (2006):
Theorem 12 (Lehmann & Casella (2006) Theorem 6.6).
Consider an experiment with parameter $\theta\in\Omega$ for a product of open sets $\Omega\in\mathbb{R}^{n}$. Assume the likelihood $l(Z,\theta)$ has the same support for all $\theta$ and finite derivatives $\frac{dl(Z,\theta)}{d\theta}$. Assume the Fisher information $\ I(\theta)=\operatorname*{\mathbb{E}}_{Z\sim l(Z,\theta)}\left[\frac{\delta^{2}\log l(Z;\theta)}{\delta\theta^{2}}\right]$ is positive definite. Then for any statistic $\delta$, we have $\operatorname*{\mathrm{Var}}_{\theta}(\delta)\geq\frac{d\operatorname*{\mathbb{E}}_{\theta}\delta}{d\theta}^{t}I(\theta)^{-1}\frac{d\operatorname*{\mathbb{E}}_{\theta}\delta}{d\theta}$. In particular, if $\operatorname*{\mathbb{E}}_{\theta}\delta$ equals the $i$-th component of $\theta$, we have
$\operatorname*{\mathrm{Var}}_{\theta}(\delta)\geq(I(\theta)^{-1})_{(i,i)}.$
We again drop the dependence on $m$ in order to declutter notation. We then use that the likelihoods of independent observations factor and the linearity of derivatives to obtain
$I(b,q,p)=nI_{n}(b,q,p)+NI_{N}(b,q,p),$
where $I(b,q,p)$ is the Fisher information for our experiment and $I_{n}(b,q,p)$ and $I_{N}(b,q,p)$ are the Fisher informations for a single $(s(m),\tilde{s}(m))$ or a single $\tilde{s}(m)$ sample, respectively. Denoting the corresponding likelihoods $l_{n}$ and $l_{N}$ respectively, we have:
$l_{n}(s=0,\tilde{s}=0)=(1-b)q$
$l_{n}(s=0,\tilde{s}=1)=(1-b)(1-q)$
$l_{n}(s=1,\tilde{s}=0)=b(1-p)$
$l_{n}(s=1,\tilde{s}=1)=bp$
$l_{N}(\tilde{s}=0)=(1-p)(1-q)+bp$
$l_{N}(\tilde{s}=1)=b(1-p)+(1-b)q.$
Computing the Fisher information [^2] We used Sympy for automatic symbolic differentiation and simplification for the remainder of the proof. The code can be found in Proofs.ipynb in the supplementary material. yields
$(I(\theta))_{(b,b)}=$ $-\frac{Nb\left(b-1\right)\left(p+q-1\right)^{2}\left(bp-b\left(p-1\right)-q\left(b-1\right)+\left(b-1\right)\left(q-1\right)\right)}{b\left(b-1\right)\left(bp+\left(b-1\right)\left(q-1\right)\right)\left(b\left(p-1\right)+q\left(b-1\right)\right)}$
$-\frac{n\left(bp+\left(b-1\right)\left(q-1\right)\right)\left(b\left(p-1\right)+q\left(b-1\right)\right)}{b\left(b-1\right)\left(bp+\left(b-1\right)\left(q-1\right)\right)\left(b\left(p-1\right)+q\left(b-1\right)\right)}$
$(I(\theta))_{(q,q)}=$ $N\left(\frac{\left(b-1\right)^{2}}{-b\left(p-1\right)-q\left(b-1\right)}+\frac{\left(b-1\right)^{2}}{bp+\left(b-1\right)\left(q-1\right)}\right)+n\left(\frac{b-1}{q-1}-\frac{b-1}{q}\right)$
$(I(\theta))_{(p,p)}=$ $-N\left(\frac{b^{2}}{b\left(p-1\right)+q\left(b-1\right)}-\frac{b^{2}}{bp+\left(b-1\right)\left(q-1\right)}\right)-n\left(\frac{b}{p-1}-\frac{b}{p}\right)$
$(I(\theta))_{(b,q)}=$ $\frac{N\left(-bp-bq+b+p+q-1\right)}{b^{2}p^{2}+2b^{2}pq-2b^{2}p+b^{2}q^{2}-2b^{2}q+b^{2}-2bpq+bp-2bq^{2}+3bq-b+q^{2}-q}$
$(I(\theta))_{(b,p)}=$ $\frac{Nb\left(-p-q+1\right)}{b^{2}p^{2}+2b^{2}pq-2b^{2}p+b^{2}q^{2}-2b^{2}q+b^{2}-2bpq+bp-2bq^{2}+3bq-b+q^{2}-q}$
$(I(\theta))_{(q,p)}=$ $\frac{Nb\left(b-1\right)\left(-bp+b\left(p-1\right)+q\left(b-1\right)-\left(b-1\right)\left(q-1\right)\right)}{\left(bp+\left(b-1\right)\left(q-1\right)\right)\left(b\left(p-1\right)+q\left(b-1\right)\right)}.$
With this, we can calculate
$(I(\theta))^{-1}_{(b,b)}$
$=\frac{b\left(Nb\left(bp-p+\left(1-b\right)\left(1-q\right)\right)^{2}\right)}{n\left(N+n\right)\left(bp+\left(b-1\right)\left(q-1\right)\right)\left(bp+\left(1-b\right)\left(1-q\right)-1\right)}$
$-\frac{b\left(\left(N+n\right)\left(b-1\right)\left(bp+\left(b-1\right)\left(q-1\right)\right)\left(bp+\left(1-b\right)\left(1-q\right)-1\right)\right)}{n\left(N+n\right)\left(bp+\left(b-1\right)\left(q-1\right)\right)\left(bp+\left(1-b\right)\left(1-q\right)-1\right)}$
$=\frac{b\left(-Nb\left(\operatorname*{\mathbb{E}}\tilde{s}-p)\right)^{2}+(1-b)\left(N+n\right)\operatorname*{\mathbb{E}}\tilde{s}(1-\operatorname*{\mathbb{E}}\tilde{s})\right)}{n\left(N+n\right)\operatorname*{\mathbb{E}}\tilde{s}(1-\operatorname*{\mathbb{E}}\tilde{s})}$
$=\frac{b(1-b)}{n}-\frac{b\left(Nb\left(\operatorname*{\mathbb{E}}\tilde{s}-p)\right)^{2}\right)}{n\left(N+n\right)\operatorname*{\mathbb{E}}\tilde{s}(1-\operatorname*{\mathbb{E}}\tilde{s})}$
$=\frac{\operatorname*{\mathrm{Var}}s}{n}-\frac{N}{n\left(N+n\right)}\frac{\left(b\operatorname*{\mathbb{E}}\tilde{s}-bp)\right)^{2}}{\operatorname*{\mathbb{E}}\tilde{s}(1-\operatorname*{\mathbb{E}}\tilde{s})}$
$=\frac{\operatorname*{\mathrm{Var}}s}{n}-\frac{1}{n+\frac{n^{2}}{N}}\frac{\operatorname*{\mathrm{Cov}}(s,\tilde{s})^{2}}{\operatorname*{\mathrm{Var}}\tilde{s}}=\operatorname*{\mathrm{Var}}\hat{\theta}^{PP}_{\lambda^{*}}.$
C.2 Theorem 6
Proof.
To proof the theorem, we reparameterize to simplify the constraints, namely setting $x=\operatorname*{\mathrm{AG}}(m)\geq 0.5$ and $\delta=b(m)-\operatorname*{\mathrm{AG}}(m)>0$. For notational convenience, we also set $q=q(m)$. We then optimize the squared correlation $\rho^{2}$ over the parameters in succession: At first, we take the derivative with respect to $q$, finding two roots. Thus, the maximum of $\rho^{2}$ with respect to $q$ is either at one of these roots or the boundary values for $q$. We consider each candidate separately, and optimize over $x$ and $\delta$ in a similar manner. This way, we enumerate all points $q,x,\delta$ that are candidates for the global maximum, i.e. all such points that are feasible, and either hit the boundary of a constraint or have zero derivative respectively, in all of the three variables.
We now delve into the details of the proof: Note that the agreement rate is given by $\operatorname*{\mathrm{AG}}(m)=(1-b)q+bq$. We therefore replace $p(m)=\frac{(b-1)q+\operatorname*{\mathrm{AG}}(m)}{b}=\frac{(\delta+x-1)q+x}{b}$ and obtain a formula for $\rho^{2}$ based on $q,x$ and $\delta:$
$\rho(s(m),\tilde{s}(m))^{2}=\frac{\left(\delta+x-1\right)\left(2\delta q-\delta+2qx-q\right)^{2}}{\left(\delta+x\right)\left(2\delta q-\delta+2qx-2q\right)\left(2\delta q-\delta+2qx-2q+1\right)}.$
Maximizing in $q\in[0,1]$ We then take the derivative with respect to $q$ to obtain
$\frac{d}{dq}\rho(s(m),\tilde{s}(m))^{2}=\frac{2\left(\delta+x-1\right)\left(2\delta q-\delta+2qx-q\right)\left(2\delta qx-\delta q-\delta x+2qx^{2}-3qx+q\right)}{\left(\delta+x\right)\left(2\delta q-\delta+2qx-2q\right)^{2}\left(2\delta q-\delta+2qx-2q+1\right)^{2}}.$
The numerator is quadratic in $q$, so it has at most two zeros. These turn out to be at $z_{1}=\frac{\delta}{2\delta+2x-1}$ and $z_{2}=\frac{\delta x}{\left(2x-1\right)\left(\delta+x-1\right)}$. We note, that there are singularities at $q=\frac{\delta}{2\left(\delta+x-1\right)}$ and $q=\frac{\delta-1}{2\left(\delta+x-1\right)}$. However, as $\delta+x<1$, the first singularity occurs at a value of $q\leq 0$. Similarly, the second singularity can be seen to occur at $q\geq 1$, as $\delta-1\leq 2\delta+2x-2$ is equivalent to $1\leq\delta+2x$, which is true as $x\geq 0.5$ and $\delta\geq 0$. Correspondingly, the singularities do not introduce any discontinuities in $q$ for $q\in(0,1)$. This means that for fixed $x$ and $\delta,$ $\rho^{2}$ is either maximized at $z_{1}$ or $z_{2}$, or at the extreme values $q=0$ or $q=1$.
Case 1:
Inserting $q=z_{1}$ turns the numerator of $\rho^{2}$ to zero, such that $\rho^{2}=0$.
Case 2:
We next analyze $z_{2}=\frac{\delta x}{\left(2x-1\right)\left(\delta+x-1\right)}$: While $\delta x$ and $12x-$ 1 are positive by assumption, $\delta+x-1$ is negative. Correspondingly, $q=z_{2}$ is not a valid value for $q$ and the maximizer must be one of the other three values.
Case 3:
Next, inserting $q=0$ yields
$\rho(s(m),\tilde{s}(m))^{2}=\frac{\delta\left(\delta+x-1\right)}{\left(\delta-1\right)\left(\delta+x\right)}.$
Maximizing in $x\in[0.5,1)$ We take the derivative with respect to $x$ to obtain
$\frac{d}{dx}\rho(s(m),\tilde{s}(m))^{2}=\frac{\delta}{\left(\delta-1\right)\left(\delta+x\right)^{2}}.$
This is negative because $\delta$ is positive while $\delta-1$ is negative. Correspondingly, it is maximized at the smallest possible value of $x=0.5$. Inserting $x=0.5$ yields
$\rho(s(m),\tilde{s}(m))^{2}=\frac{\delta\left(\delta-0.5\right)}{\delta^{2}-0.5\delta-0.5}.$
Maximizing in $\delta\in(0,1-x)$ We take the derivative with respect to $\delta$, obtaining
$\frac{d}{d\delta}\rho(s(m),\tilde{s}(m))^{2}=-\frac{\left(\delta-0.25\right)}{\left(\delta-1\right)^{2}\left(\delta+0.5\right)^{2}}.$
This is zero at $\delta=0.25$ and negative for larger $\delta$, indicating a local maximum. We thus insert $\delta=0.25$, obtaining
$\rho(s(m),\tilde{s}(m))^{2}=\frac{0.25\left(0.25-0.5\right)}{0.0625-0.125-0.5}=\frac{1}{9}.$
Case 4:
This leaves us with $q=1$. We again insert, obtaining
$\rho(s(m),\tilde{s}(m))^{2}=\frac{\left(\delta+x-1\right)\left(\delta+2x-1\right)}{\left(\delta+x\right)\left(\delta+2x-2\right)}.$
Maximizing in $\delta\in(0,1-x)$ Taking the derivative with respect to delta yields
$\frac{d}{d\delta}\rho(s(m),\tilde{s}(m))^{2}=\frac{\left(x-1\right)\left(2\delta+3x-2\right)}{\left(\delta+x\right)^{2}\left(\delta+2x-2\right)^{2}}.$
Note that this only has singularities at $\delta+x=1$ and $\delta+2x=2$, both of which are ruled out by the constraint $\delta+x<1$. The derivative is zero at $\delta=1-\frac{3x}{2}$ and negative for larger $\delta$, indicating a local maximum. When $x>\frac{2}{3}$, this value violates the constraint of $\delta>0$, such that the maximum has to be at $\delta=0$ or $\delta=1-x$ instead.
Local maximizing in $x\in(0,\frac{2}{3})$ We insert the local maximizer $\delta=1-\frac{3x}{2}$,
$\rho(s(m),\tilde{s}(m))^{2}=\frac{x^{2}}{\left(x-2\right)^{2}}$
with positive derivative
$\frac{d}{dx}\rho(s(m),\tilde{s}(m))^{2}=-\frac{4x}{\left(x-2\right)^{3}}.$
This means that the maximum value is at $x=\frac{2}{3}$, where
$\rho(s(m),\tilde{s}(m))^{2}=\frac{1}{4}.$
Boundary maximizing in $x\in(0,1)$ We check the boundaries for $\delta$, i.e., $0<\delta<1-x$. First, we notice that for $\delta\to 1-x$, $\rho^{2}\to 0$. We thus focus on $\delta\to 0$. Inserting yields
$\rho(s(m),\tilde{s}(m))^{2}=\frac{2x-1}{2x}$
with positive derivative
$\frac{d}{dx}\rho(s(m),\tilde{s}(m))^{2}=\frac{1}{2x^{2}}.$
Correspondingly, it is maximized for $x\to 1$, where it clearly goes to $0.5$. With this, the maximum attainable value for $\rho^{2}$ equals $0.5$.
C.3 Theorem 9
Proof.
We again drop the dependence on $m$ for notational convenience.
$\rho(s(m),\tilde{s}(m))^{2}$
$=b\frac{(p-((1-q)(1-b)+pb))^{2}}{((1-q)(1-b)+pb)(1-((1-q)(1-b)+pb))(1-b)}$
$=b\frac{((1-b)(p+q-1))^{2}}{(1-q-b+qb+pb)(1-(1-q-b+qb+pb))(1-b)}$
$=b(1-b)\frac{(2\operatorname*{\mathrm{BA}}-1)^{2}}{(1-q-b+2b\operatorname*{\mathrm{BA}})(q+b-2b\operatorname*{\mathrm{BA}})}.$
Clearly, this equals zero when $\operatorname*{\mathrm{BA}}=0.5$. Fixing $\operatorname*{\mathrm{BA}}\neq 0.5$ and taking the derivative with respect to $q$ yields
$\frac{d}{dq}\rho(s(m),\tilde{s}(m))^{2}=-\frac{b(1-b)(2\operatorname*{\mathrm{BA}}-1)^{2}(4b\operatorname*{\mathrm{BA}}-2b-2q+1)}{(1-q-b+2b\operatorname*{\mathrm{BA}})^{2}(q+b-2b\operatorname*{\mathrm{BA}})^{2}}.$
Assuming $10<b<$ 1 and $\operatorname*{\mathrm{BA}}\neq 0.5$, this is zero if and only if $-(4b\operatorname*{\mathrm{BA}}-2b-2q+1)$ is zero and clearly positive for larger $q$. Correspondingly, as long as the singularities in $\rho^{2}$ lie outside of the $q-$domain, $q=2b\operatorname*{\mathrm{BA}}-b+\frac{1}{2}$ is a unique minimum, while $\rho(s(m),\tilde{s}(m))^{2}$ is maximized either at the minimal or the maximal possible value of $q$.
We first insert the minimum, noting that for a lower bound we do not need to worry about whether this value $q$ is indeed attainable. This yields
$\rho(s(m),\tilde{s}(m))^{2}=4b(1-b)(2\operatorname*{\mathrm{BA}}-1)^{2},$
as both denominator terms reduce to $\frac{1}{2}.$
Next, we assume $\operatorname*{\mathrm{BA}}>\frac{1}{2}$. As both $p$ and $q$ are in $(0,1)$, we get that $q=2\operatorname*{\mathrm{BA}}-p\geq 2\operatorname*{\mathrm{BA}}-1$. Now, the singularities occur at $q=b(2\operatorname*{\mathrm{BA}}-1)$ and $q=2\operatorname*{\mathrm{BA}}b-b+1$, the first of which is smaller than $12\operatorname*{\mathrm{BA}}-$ 1, while the second is larger than $1$. Correspondingly, we can ignore them and insert the minimal $q=2\operatorname*{\mathrm{BA}}-1$. We obtain
$\rho(s(m),\tilde{s}(m))^{2}=\frac{b(2\operatorname*{\mathrm{BA}}-1)}{2b\operatorname*{\mathrm{BA}}-2\operatorname*{\mathrm{BA}}-b+2},$
and inserting $q=1$, we obtain
$\rho(s(m),\tilde{s}(m))^{2}=\frac{(b-1)(2\operatorname*{\mathrm{BA}}-1)}{2b\operatorname*{\mathrm{BA}}-b-1}.$
These two terms are symmetric in the sense that replacing $b$ in the first term with $1-b$ yields the second term and vice versa. Thus, it is sufficient to maximize the second term with respect to $b$. We obtain
$\frac{d}{db}\rho(s(m),\tilde{s}(m))^{2}=\frac{(2\operatorname*{\mathrm{BA}}-2)(2\operatorname*{\mathrm{BA}}-1)}{(2b\operatorname*{\mathrm{BA}}-b-1)^{2}},$
which is negative such that $\rho(s(m),\tilde{s}(m))^{2}$ is maximized at $b=0$, where it equals $12\operatorname*{\mathrm{BA}}-$ 1.
Similarly, for $\operatorname*{\mathrm{BA}}<\frac{1}{2}$, $q$ can range from 00 to $2\operatorname*{\mathrm{BA}}$. Again, the singularities are at $q=b(2\operatorname*{\mathrm{BA}}-1)$ and $q=2\operatorname*{\mathrm{BA}}b-b+1$. This time, the first one is negative, while the second one can easily be seen to be larger than $2\operatorname*{\mathrm{BA}}$. This again allows us to ignore the singularities. Inserting the extreme values yields $\frac{(2\operatorname*{\mathrm{BA}}-1)(b-1)}{2b\operatorname*{\mathrm{BA}}-b+1}$ and $\frac{b(2\operatorname*{\mathrm{BA}}-1)}{2b\operatorname*{\mathrm{BA}}-2\operatorname*{\mathrm{BA}}-b}$, which again are the same after replacing $b$ with $1-b$. Taking the derivative of the first term with respect to $b$ yields
$\frac{d}{db}\rho(s(m),\tilde{s}(m))^{2}=\frac{(2\operatorname*{\mathrm{BA}})(2\operatorname*{\mathrm{BA}}-1)}{(2b\operatorname*{\mathrm{BA}}-b+1)^{2}},$
which is negative as $\operatorname*{\mathrm{BA}}<\frac{1}{2}$. Correspondingly, $b=0$ now maximizes $\rho(s(m),\tilde{s}(m))^{2}$, at which point it equals $1-2\operatorname*{\mathrm{BA}}$. Joining both cases, we thus obtain
$\rho(s(m),\tilde{s}(m))^{2}\leq|2\operatorname*{\mathrm{BA}}-1|.$
C.4 Proposition 1
Proof.
$\operatorname*{\mathbb{E}}\tilde{s}(x,m_{i})$ $=\operatorname*{\mathbb{P}}(\tilde{m}(x)=m_{i}(x))$
$=\operatorname*{\mathbb{P}}(\tilde{m}(x)=y(x))\operatorname*{\mathbb{P}}(\tilde{m}(x)=m_{i}(x)|\tilde{m}(x)=y(x))+\operatorname*{\mathbb{P}}(m(x)=m_{i}(x),\tilde{m}(x)\neq y(x))$
$=\operatorname*{\mathbb{E}}s(\tilde{m})\cdot 1+\operatorname*{\mathbb{P}}(m_{i}(x)\neq y(x),\tilde{m}(x)\neq y(x))$
$=\operatorname*{\mathbb{E}}s(\tilde{m})+\operatorname*{\mathbb{P}}(m_{i}(x)\neq y(x))$
$=\operatorname*{\mathbb{E}}s(\tilde{m})+1-\operatorname*{\mathbb{E}}s(m_{j}),$
which is monotonously falling in $\operatorname*{\mathbb{E}}s(m_{j})$. ∎
C.5 Proposition 2
Proof.
We again drop the dependence on $m$. In the first case, we need that
$bp+(1-b)q=\operatorname*{\mathrm{AG}}=r=1-\operatorname*{\mathrm{JB}}=1-(1-q)(1-b)+(1-p)b.$
This is equivalent to
$bp+q-bq=q+b-qb+b-bp$
or
$bp=b+b-bp,$
i.e. $p=1$. With this, the constraint of $\operatorname*{\mathrm{AG}}=r$ turns into $b+q(1-b)=r$. The left side is clearly continuous and monotonous in $q$ and inserting the extreme values of zero and one, we obtain $r=b$ and $r=1$, such that we can find a solution for all intermediate values of $r$.
In the second case, we want
$bp+(1-b)q=\operatorname*{\mathrm{AG}}=r=\operatorname*{\mathrm{JB}}+1=(1-q)(1-b)-(1-p)b+1.$
We simplify to get
$bp+q-bq=1-q-b+qb-b+pb+1$
or equivalently
$q-bq=-q-b+qb-b+2$
or
$2q-2bq+2b=2,$
which is achieved at $q=1$. In that case, the constraint of $\operatorname*{\mathrm{AG}}=r$ turns into $bp+(1-b)=r$ with extreme values again lying at $1-b=r$ and $1=r$. ∎
C.6 Proposition 4
Proof.
We first note that $\operatorname*{\mathrm{Var}}s(m)=b(m)(1-b(m))$ and $\operatorname*{\mathrm{Var}}\tilde{s}(m)=\operatorname*{\mathbb{E}}\tilde{s}(m)(1-\operatorname*{\mathbb{E}}\tilde{s}(m)),$ as both $s$ and $\tilde{s}$ are binary. Furthermore, $\operatorname*{\mathbb{E}}[s(m)\tilde{s}(m)]=b(m)p(m)$, such that $\operatorname*{\mathrm{Cov}}(s(m),\tilde{s}(m))=b(m)(p(m)-\operatorname*{\mathbb{E}}\tilde{s}(m))$. With this, it is easy to see that the formula for $\rho^{2}$ in the proposition statement is indeed correct. With this, we take the limit of $N\to\infty$ to maximize the variance reduction and plug in our parameterization for $s$ and $\tilde{s}$ to simplify
$1-\frac{\operatorname*{\mathrm{Var}}\hat{\theta}^{PP}_{\lambda^{*}}}{\operatorname*{\mathrm{Var}}\hat{\theta}}=\rho(s(m),\tilde{s}(m))^{2}$ $=\frac{b(m)}{(1-b(m))}\frac{(p(m)-\operatorname*{\mathbb{E}}\tilde{s}(m))^{2}}{\operatorname*{\mathbb{E}}\tilde{s}(m)(1-\operatorname*{\mathbb{E}}\tilde{s}(m))}\,.\qed$